# Dynamic Scene Understanding through Object-Centric Voxelization and Neural Rendering

### v1

@@ 1 Introduction
We build the model, DynaVol-S, in an inverse graphics framework based on neural radiance field (NeRF) to effectively leverage the multi-view video frames collected from a monocular camera. However, the classic canonical-space rendering pipeline does not support object-centric learning. We identify two key challenges: First, how to decouple the underlying dynamics of each object from its visual appearance and capture the local geometric structures of each object. Second, how to integrate the global semantic priors of the scene, typically extracted from 2D foundation models, into 3D volume rendering, which is crucial for inferring 3D geometries given partial observations, complex real-world structures, and severe occlusions.
@@
For the second challenge of semantics integration, we observe that extending our previous work of DynaVol [10] to real-world data poses significant challenges due to the increased complexity of color and geometry patterns and the limited range of available view directions. To address these issues, we present a new method that leverages 2D semantic feature maps extracted from an image encoding network trained on large-scale datasets. We develop a novel semantic volume slot attention module to re-project the 2D feature maps onto 3D voxel grids. This design enables our model to correlate the object-centric voxel grids with the pre-learned 2D feature maps, incorporating common knowledge of object geometries. This is particularly beneficial for handling previously unseen object structures, enhancing the model’s ability to interpret complex real-world scenes.
@@ 3.1 Preliminaries on NeRF
The outputs $\sigma$ and $\mathbf{c}$ are then transformed to the pixel values on a 2D image using volume rendering techniques [18]. Specifically, given a ray $\mathbf{r}=\mathbf{o}+t\mathbf{d}$ emitted from $\mathbf{x}$ to a specific pixel in the image, where $\mathbf{o}$ and $\mathbf{d}$ are respectively the origin and direction of the ray and $t$ is the distance from $\mathbf{x}$ to the camera. The color of the pixel is determined by the integral of the $(\sigma_{i},\mathbf{c}_{i})$ values of all the points along the ray within a predefined near bound $t_{n}$ and a far bound $t_{f}$:
$\begin{split}C(\mathbf{r})=&\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathbf{c}_{i},\\ T_{i}=&\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right),\end{split}$ (2)
where $P$ denotes the number of sampled points, and $\delta_{i}$ denotes the distance between the $i$-th and $(i+1)$-th points. In this work, we learn object-centric representations that evolve over time to drive a compositional NeRF renderer.
Figure 2: DynaVol-S consists of three groups of network components: the bi-directional dynamics modules ($f_{\psi}$, $f_{\xi}^{\prime}$), the volume slot attention module, and neural rendering modules based on 3D and 4D voxels respectively. The training scheme involves a warmup stage (top), an object-centric voxel grids initialization stage (middle, whose pseudocode is given in Alg. 1), and a multi-grids joint optimization stage (bottom). For better clarity, we present the symbols and their corresponding descriptions in Table I.
@@
$\mathcal{V}^{\text{M}}$ Voxel grids for occupancy probabilities per object $\widehat{C}$ Rendered 2D pixel colors
$\mathcal{V}^{\text{D}}$ Voxel grids for opacity $C$ Observed 2D pixel colors
$\mathcal{V}^{\text{C}}$ Voxel grids for color-related features $\mathcal{A}$ Semantic probabilities
$\mathcal{V}^{\text{S}}$ Voxel grids for semantic features $\widehat{\mathcal{A}}$ Rendered 2D semantic probabilities
@@
$\mathbf{S}$ / $\{\mathbf{s}_{n}\}_{n=1}^{N}$ Object-centric slot features $N_{\phi},N_{\phi^{\prime}}$ Compositional/Non-compositional NeRF
@@ 3.2 Object-Centric Voxelization
DynaVol-S is trained in an inverse graphics framework to synthesize $\{\mathbf{I}_{t}\}_{t=1}^{T}$ without any further supervision. Formally, the goal is to learn an object-centric projection of $(\mathbf{x},\mathbf{d},t,\{\mathbf{z}_{n}\}_{n=1}^{N},\mathcal{V}^{\text{M}},\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}})\xrightarrow[]{}(\mathbf{c},\sigma)$. $\mathbf{x}=(x,y,z)$ is a 3D point, $\{\mathbf{z}_{n}\}_{n=1}^{N}\in\mathbb{R}^{N\times D_{z}}$ denotes a set of learnable latent codes associated with individual objects, and $N$ is the predefined number of the latent codes that is assumed to be larger than the number of objects. $\mathcal{V}^{\text{M}},\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}}$ are voxelized features, which respectively represent occupancy probabilities per object, opacity, and color-related features.
To understand the scene from an object-centric perspective, we extend the classic approach of using 3D voxel grids for volume rendering by incorporating 4D voxel grids, represented as $\mathcal{V}^{\text{M}}\in\mathbb{R}^{N\times N_{x}\times N_{y}\times N_{z}}$, where $(N_{x},N_{y},N_{z})$ denote the spatial resolutions. The newly added dimension in the expanded voxel features capture the occupancy probabilities for each object at the grid vertices. Therefore, the occupancy probabilities $\{p_{n}\}_{n=1}^{N}$ at any given 3D point within the grid cells are efficiently retrieved using trilinear interpolation:
$\mathrm{Interp}(\mathbf{x},\mathcal{V}^{\text{M}}):\ (\mathbb{R}^{3},\mathbb{R}^{N\times N_{x}\times N_{y}\times N_{z}})\xrightarrow{}\mathbb{R}^{N}.$ (3)
Following this, we apply the Softmax activation function to the interpolated values to normalize $\{p_{n}\}_{n=1}^{N}$, ensuring they sum to one and are proportionately scaled.
In addition to $\mathcal{V}^{\text{M}}$, our approach employs three other voxelized representations that are independent of the object index: $\mathcal{V}^{\text{D}}\in\mathbb{R}^{1\times N_{x}\times N_{y}\times N_{z}}$ for modeling the opacity $\sigma$, $\mathcal{V}^{\text{C}}\in\mathbb{R}^{D_{c}\times N_{x}\times N_{y}\times N_{z}}$ for modeling color-related features $\mathbf{f}^{c}$, and $\mathcal{V}^{\text{S}}\in\mathbb{R}^{D_{s}\times N_{x}\times N_{y}\times N_{z}}$ for modeling semantic features of the scene. The incorporation of semantic features is a pivotal contribution of the improved DynaVol-S. We optimize the initial states of $\{\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{S}},\mathcal{V}^{\text{M}}\}$ and learn a deformation module $f_{\psi}(\cdot)$ to transform them over time in the canonical space.
@@ 3.3 Model Components
    The semantic volume slot attention module, which learn the semantic features $\mathcal{V}^{\text{S}}$ and disentangle them into object-centric slot features $\mathbf{S}$ respectively. We extract a set of 2D semantic-level features $\{\mathbf{F}_{t}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}\}_{t=1}^{T}$ from $\{\mathbf{I}_{t}\}_{t=1}^{T}$ using a pre-trained 2D feature extractor, e.g., DINOv2 [19] as additional learning objectives for this module.
@@
    The neural renders $N_{\phi}$ and $N_{\phi^{\prime}}$. $N_{\phi}$ is a compositional NeRF that jointly uses $\mathcal{V}^{\text{D}},\mathcal{V}^{\text{C}},\mathcal{V}^{\text{M}}$, and $\{\mathbf{z}_{n}\}_{n=1}^{N}$ for rendering of each object. $N_{\phi^{\prime}}$ is a non-compositional NeRF only used to initialize the voxel representations.
@@ 3.3.1 Bi-Directional Deformation Networks
$\small\begin{split}\sigma_{i}=\mathrm{Interp}&\left(\mathbf{x}_{i}^{\prime},\mathcal{V}^{\text{D}}\right),\\ \mathbf{f}_{i}^{c}=\mathrm{Interp}\left(\mathbf{x}_{i}^{\prime},\mathcal{V}^{\text{C}}\right)&\oplus\ldots\oplus\mathrm{Interp}\left(\mathbf{x}_{i}^{\prime},\mathcal{V}^{\text{C}}\left[::s_{M}\right]\right).\end{split}$ (4)
@@
Forward dynamics modeling. Additionally, we use another dynamics module $f_{\xi}^{\prime}(\mathbf{x}_{i},t)$ to model the forward movement $\Delta x_{i}^{\prime}$ from the initial time to timestamp $t$. This module enables the calculation of a cycle-consistency loss, enhancing the coherence of the learned canonical-space transitions. It provides useful information representing the forward dynamics for the subsequent connect components algorithm in the warmup stage (described in detail in Sec. 3.4), which can improve the initialization of the aforementioned voxel grids.
@@ 3.3.2 Semantic Volume Slot Attention
Volume slot attention. During the warmup phase, we employ the volume slot attention to progressively derive a set of global time-invariant object-level representations from the local, time-varying voxel features. Specifically, we introduce the semantic voxel grids, $\mathcal{V}^{\text{S}}\in\mathbb{R}^{D_{s}\times N_{x}\times N_{y}\times N_{z}}$, and a set object-centric “slots features”, $\mathbf{S}\in\mathbb{R}^{N\times D_{s}}$, borrowed from the prior literature on 2D static scene decomposition [21]. Given a 3D point $\mathbf{x}$ at timestamp $t$, we have:
$\begin{split}\mathbf{f}_{i}^{s}\in\mathbb{R}^{D_{s}}&=\text{Interp}\left(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),\mathcal{V}^{\text{S}}\right)\\ \mathcal{A}_{i}\in\mathbb{R}^{N}&=\mathrm{softmax}_{N}\left(\frac{\mathbf{f}_{i}^{s}\cdot\mathbf{S}^{T}}{\tau}\right)\\ \mathbf{f}_{i}^{s}\in\mathbb{R}^{D_{s}}&=\mathcal{A}_{i}\cdot\mathbf{S},\end{split}$ (5)
where $T$ denotes the transpose of a matrix, $\tau$ is a learnable temperature factor, and $\mathcal{A}_{i}$ represents the semantic probabilities that are further used to initialize the 4D voxel grids, $\mathcal{V}^{\text{M}}$. Unlike prior slot attention methodologies, we do not use query, key, and value projections. This approach improves efficiency and ensures that linear projections do not affect the correlation between features and slots. Subsequently, $\mathbf{f}^{s}_{i}$ and $\mathcal{A}_{i}$ are used to obtain the 2D semantic features and corresponding 2D maps of semantic probabilities with the quadrature rule [18]:
@@ 3.3.3 Object-Centric Renderer
$\begin{split}p_{i}=&\ \mathrm{softmax}(\mathrm{Interp}(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),\mathcal{V}^{\text{M}});\mathrm{Dim}=N),\\ \overline{\mathbf{z}}_{i}=&\sum_{n=1}^{N}p_{i}[n]\cdot\mathbf{z}_{n}.\end{split}$ (7)
Subsequently, we use an MLP to learn object-centric projections $N_{\phi}:(\mathbf{f}_{i}^{c},\mathbf{d},\overline{\mathbf{z}}_{i})\xrightarrow{}\mathbf{c}_{i}$, where $\mathbf{f}_{i}^{c}$ are color-related features from Eq. (4). We apply positional encodings to $\mathbf{f}_{i}^{c}$ and $\mathbf{d}$ before feeding them to the MLP. In the rendering process, we directly query the opacity $\sigma$ from the voxel grids $\mathcal{V}^{\text{D}}$ within the canonical space. We follow Eq. (2) to estimate the pixel color of a sampling ray using the predicted $\mathbf{c}_{i}$ and the retrieved $\sigma_{i}$.

### v2

@@ 1 Introduction
We propose DynaVol-S, an inverse graphics model based on neural radiance field (NeRF), to learn 3D scene layouts from sequential video frames captured by a moving monocular camera. However, the classic canonical-space rendering pipeline does not support object-centric learning. We identify two key challenges: First, how to decouple the underlying dynamics of each object from its visual appearance and capture the local geometric structures of each object. Second, how to integrate the global semantic priors of the scene, typically extracted from 2D foundation models, into 3D volume rendering, which is crucial for inferring 3D geometries given partial observations, complex real-world structures, and severe occlusions.
@@
For the second challenge of semantics integration, we observe that extending our previous work of DynaVol [10] to real-world data poses significant challenges due to the increased complexity of color and geometry patterns and the limited range of available view directions. To address these issues, we present a new method that leverages 2D semantic feature maps extracted from an image encoding network trained on large-scale datasets. We develop a novel semantic volume slot attention module to project the learned 3D voxel grids onto 2D feature maps, This design enables our model to correlate the object-centric voxel grids with the pre-learned 2D feature maps, incorporating common knowledge of object geometries. This is particularly beneficial for handling previously unseen object structures, enhancing the model’s ability to interpret complex real-world scenes.
@@ 3.1 Preliminaries on NeRF
The outputs $\sigma$ and $\mathbf{c}$ are then transformed to the pixel values on a 2D image using volume rendering techniques [18]. Given a ray $\mathbf{r}=\mathbf{o}+t\mathbf{d}$ emitted from $\mathbf{x}$ to a specific pixel in the image, where $\mathbf{o}$ and $\mathbf{d}$ are respectively the origin and direction of the ray and $t$ is the distance from $\mathbf{x}$ to the camera. The color of the pixel is determined by the integral of $(\sigma_{i},\mathbf{c}_{i})$ of all the points along the ray within a predefined near bound $t_{n}$ and a far bound $t_{f}$, where $P$ denotes the number of sampled points, and $\delta_{i}$ denotes the distance between the $i$-th and $(i+1)$-th points:
$\begin{split}C(\mathbf{r})=&\sum_{i=1}^{P}T_{i}\left(1-\exp(-\sigma_{i}\delta_{i})\right)\mathbf{c}_{i},\\ T_{i}=&\exp\left(-\sum_{j=1}^{i-1}\sigma_{j}\delta_{j}\right).\end{split}$ (2)
In this work, we learn object-centric representations that evolve over time to drive a compositional NeRF renderer.
Figure 2: DynaVol-S consists of three groups of network components: the bi-directional dynamics modules ($f_{\psi}$, $f_{\xi}^{\prime}$), the volume slot attention module, and neural rendering modules based on 3D and 4D voxels respectively. The training scheme involves a warmup stage (top), an object-centric voxel grids initialization stage whose pseudocode is given in Alg. 1, and a multi-grids joint optimization stage (bottom). For clarity, we denote $\mathcal{V}^{\text{Color}}$ and $\mathcal{V}^{\text{Opac}}$ collectively as $\mathcal{V}^{\text{C/O}}$ and present the symbols and their descriptions in Table I.
@@
$\mathcal{V}^{\text{Occ}}$ Voxel grids for occupancy probabilities per object $\widehat{C}$ Rendered 2D pixel colors
$\mathcal{V}^{\text{Opac}}$ Voxel grids for opacity $C$ Observed 2D pixel colors
$\mathcal{V}^{\text{Color}}$ Voxel grids for color-related features $\mathcal{A}$ Semantic probabilities
$\mathcal{V}^{\text{Sem}}$ Voxel grids for semantic features $\widehat{\mathcal{A}}$ Rendered 2D semantic probabilities
@@
$\mathbf{S}$ / $\{\mathbf{s}_{n}\}_{n=1}^{N}$ Object-centric slot features $\mathcal{N}_{\phi},\mathcal{N}_{\phi^{\prime}}$ Compositional/Non-compositional NeRF
@@ 3.2 Object-Centric Voxelization
DynaVol-S is trained in an inverse graphics framework to synthesize $\{\mathbf{I}_{t}\}_{t=1}^{T}$ without any further supervision. Formally, the goal is to learn an object-centric projection of $(\mathbf{x},\mathbf{d},t,\{\mathbf{z}_{n}\}_{n=1}^{N},{\mathcal{V}^{\text{Occ}}},{\mathcal{V}^{\text{Opac}}},{\mathcal{V}^{\text{Color}}})\xrightarrow[]{}(\mathbf{c},\sigma)$. $\mathbf{x}=(x,y,z)$ is a 3D point, $\{\mathbf{z}_{n}\}_{n=1}^{N}\in\mathbb{R}^{N\times D_{z}}$ denotes a set of learnable latent codes associated with individual objects, and $N$ is the predefined number of the latent codes that is assumed to be larger than the number of objects. ${\mathcal{V}^{\text{Occ}}},{\mathcal{V}^{\text{Opac}}},{\mathcal{V}^{\text{Color}}}$ are voxelized features, which respectively represent occupancy probabilities per object, opacity, and color-related features.
To understand the scene from an object-centric perspective, we extend the classic approach of using 3D voxel grids for volume rendering by incorporating 4D voxel grids, represented as ${\mathcal{V}^{\text{Occ}}}\in\mathbb{R}^{N\times N_{x}\times N_{y}\times N_{z}}$, where $(N_{x},N_{y},N_{z})$ denote the spatial resolutions. The newly added dimension in the expanded voxel features captures the occupancy probabilities for each object at the grid vertices. We have
$\mathrm{Interp}(\mathbf{x},{\mathcal{V}^{\text{Occ}}}):\ (\mathbb{R}^{3},\mathbb{R}^{N\times N_{x}\times N_{y}\times N_{z}})\xrightarrow{}\mathbb{R}^{N},$ (3)
which means the occupancy probabilities $\{p_{n}\}_{n=1}^{N}$ at any given 3D position $\mathbf{x}$ within the voxel grid are efficiently retrieved using trilinear interpolation. Following this, we apply the Softmax activation function to the interpolated values to normalize $\{p_{n}\}_{n=1}^{N}$, ensuring they sum to one and are proportionately scaled.
In addition to ${\mathcal{V}^{\text{Occ}}}$, our approach employs three other voxelized representations that are independent of the object index: ${\mathcal{V}^{\text{Opac}}}\in\mathbb{R}^{1\times N_{x}\times N_{y}\times N_{z}}$ for modeling the opacity $\sigma$, ${\mathcal{V}^{\text{Color}}}\in\mathbb{R}^{D_{c}\times N_{x}\times N_{y}\times N_{z}}$ for modeling color-related features $\mathbf{f}^{c}$, and ${\mathcal{V}^{\text{Sem}}}\in\mathbb{R}^{D_{s}\times N_{x}\times N_{y}\times N_{z}}$ for modeling semantic features of the scene. The incorporation of semantic features is a pivotal contribution of the improved DynaVol-S. We optimize the initial states of $\{{\mathcal{V}^{\text{Opac}}},{\mathcal{V}^{\text{Color}}},{\mathcal{V}^{\text{Sem}}},{\mathcal{V}^{\text{Occ}}}\}$ and learn a deformation module $f_{\psi}(\cdot)$ to transform them over time in the canonical space.
@@ 3.3 Model Components
    The semantic volume slot attention module, which learn the semantic features ${\mathcal{V}^{\text{Sem}}}$ and disentangle them into object-centric slot features $\mathbf{S}$ respectively. We extract a set of 2D semantic-level features $\{\mathbf{F}_{t}\in\mathbb{R}^{C\times H^{\prime}\times W^{\prime}}\}_{t=1}^{T}$ from $\{\mathbf{I}_{t}\}_{t=1}^{T}$ using a pre-trained 2D feature extractor, e.g., DINOv2 [19] as additional learning objectives for this module.
@@
    The neural renders $\mathcal{N}_{\phi}$ and $\mathcal{N}_{\phi^{\prime}}$. $\mathcal{N}_{\phi}$ is a compositional NeRF that jointly uses ${\mathcal{V}^{\text{Opac}}},{\mathcal{V}^{\text{Color}}},{\mathcal{V}^{\text{Occ}}}$, and $\{\mathbf{z}_{n}\}_{n=1}^{N}$ for rendering of each object. $\mathcal{N}_{\phi^{\prime}}$ is a non-compositional NeRF only used to initialize the voxel representations.
@@ 3.3.1 Bi-Directional Deformation Networks
$\small\begin{split}\sigma_{i}=\mathrm{Interp}&\left(\mathbf{x}_{i}^{\prime},{\mathcal{V}^{\text{Opac}}}\right),\\ \mathbf{f}_{i}^{c}=\mathrm{Interp}\left(\mathbf{x}_{i}^{\prime},{\mathcal{V}^{\text{Color}}}\right)&\oplus\ldots\oplus\mathrm{Interp}\left(\mathbf{x}_{i}^{\prime},{\mathcal{V}^{\text{Color}}}\left[::s_{M}\right]\right).\end{split}$ (4)
@@
Forward dynamics modeling. Additionally, we employ another dynamics module $f_{\xi}^{\prime}(\mathbf{x}_{i},t)$ to capture the forward movement $\Delta x_{i}^{\prime}$ from the initial time to timestamp $t$. This module enables the calculation of a cycle-consistency loss, as defined in Eq. (9), which has been shown in previous literature [20, 13] to improve the coherence of learned canonical-space transitions. We further discuss the validity of the cycle-consistency training approach in Sec. 3.4.1. As a result, $f_{\xi}^{\prime}(\mathbf{x}_{i},t)$ provides useful information, i.e., the positions and velocities of each voxel at time $t$, for the subsequent connect components algorithm in the warmup stage (detailed in Sec. 3.4.1), which can improve the initialization of the voxel grids.
@@ 3.3.2 Semantic Volume Slot Attention
Volume slot attention. During the warmup phase, we employ the volume slot attention to progressively derive a set of global time-invariant object-level representations from the local, time-varying voxel features. Specifically, we introduce the semantic voxel grids, ${\mathcal{V}^{\text{Sem}}}\in\mathbb{R}^{D_{s}\times N_{x}\times N_{y}\times N_{z}}$, and a set object-centric “slots features”, $\mathbf{S}\in\mathbb{R}^{N\times D_{s}}$, borrowed from the prior literature on 2D static scene decomposition [22]. Given a 3D point $\mathbf{x}$ at timestamp $t$, we have:
$\begin{split}{\mathbf{f}_{i}^{s_{0}}}\in\mathbb{R}^{D_{s}}&=\text{Interp}\left(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),{\mathcal{V}^{\text{Sem}}}\right)\\ \mathcal{A}_{i}\in\mathbb{R}^{N}&=\mathrm{softmax}_{N}\left(\frac{{\mathbf{f}_{i}^{s_{0}}}\cdot\mathbf{S}^{T}}{\tau}\right)\\ \mathbf{f}_{i}^{s}\in\mathbb{R}^{D_{s}}&=\mathcal{A}_{i}\cdot\mathbf{S},\end{split}$ (5)
where $T$ denotes the transpose of a matrix, $\tau$ is a learnable temperature factor, and $\mathcal{A}_{i}$ represents the semantic probabilities that are further used to initialize the 4D voxel grids, ${\mathcal{V}^{\text{Occ}}}$. Unlike prior slot attention methodologies, we do not use query, key, and value projections. This approach improves efficiency and ensures that linear projections do not affect the correlation between features and slots. Subsequently, $\mathbf{f}^{s}_{i}$ and $\mathcal{A}_{i}$ are used to obtain the 2D semantic features and corresponding 2D maps of semantic probabilities with the quadrature rule [18]:
@@ 3.3.3 Object-Centric Renderer
$\begin{split}{p_{i}=}&\ {\mathrm{softmax}_{N}(\mathrm{Interp}(\mathbf{x}_{i}+f_{\psi}(\mathbf{x}_{i},t),{\mathcal{V}^{\text{Occ}}}),}\\ \overline{\mathbf{z}}_{i}=&\sum_{n=1}^{N}p_{i}[n]\cdot\mathbf{z}_{n}.\end{split}$ (7)
Subsequently, we use an MLP to learn object-centric projections $\mathcal{N}_{\phi}:(\mathbf{f}_{i}^{c},\mathbf{d},\overline{\mathbf{z}}_{i})\xrightarrow{}\mathbf{c}_{i}$, where $\mathbf{f}_{i}^{c}$ are color-related features from Eq. (4). We apply positional encodings to $\mathbf{f}_{i}^{c}$ and $\mathbf{d}$ before feeding them to the MLP. In the rendering process, we directly query the opacity $\sigma$ from the voxel grids ${\mathcal{V}^{\text{Opac}}}$ within the canonical space. We follow Eq. (2) to estimate the pixel color of a sampling ray using the predicted $\mathbf{c}_{i}$ and the retrieved $\sigma_{i}$.
In addition to the object-centric renderer, we also introduce a non-compositional NeRF, denoted by $\mathcal{N}_{\phi^{\prime}}$. Unlike $\mathcal{N}_{\phi}$, which is used in the multi-grids joint optimization stage, $\mathcal{N}_{\phi^{\prime}}$ is used in the warmup stage to initialize the voxel representations. Specifically, it employs an MLP to learn a mapping function: $(\mathbf{f}_{i}^{c},\mathbf{d})\xrightarrow{}\mathbf{c}_{i}$, which is unrelated to any object-centric features.